OpenStudy (doggiebounce):

Solve log(6x + 14) = 1

OpenStudy (raden):

hint : setting 1 = log 10 then cancel out the log to both sides

OpenStudy (raden):

log(6x + 14) = 1 log(6x + 14) = log 10 cancel out the log 6x+14=10 now, solve for x

OpenStudy (doggiebounce):

OK thank you so much. :)

OpenStudy (raden):

yw